OpenStudy (anonymous):

Solve using substitution x/3+y/5=1 x+6y=12

OpenStudy (anonymous):

You can make that first equation a little more palatable to work with by multiplying it by 15.

OpenStudy (anonymous):

Then if you multiply it by 2 and subtract it from the second equation, you can get y to eliminate and you'll have a solution for x.

OpenStudy (anonymous):

Hi Cliff, quick question, when you solve it in the manner you suggest, is that solve by elimination?

OpenStudy (anonymous):

i need to solve by substitution!

OpenStudy (anonymous):

@hugo, yes, I suppose it is.

OpenStudy (anonymous):

In that case, I still recommend multiplying the first equation by 15 to clear the fractions, then solve the first equation for y, and the second equation for x, then do a substitution.

OpenStudy (anonymous):

so if i do heres what i did so far \[ \frac{ x }{ 3 }+\frac{ y }{ 5} = 1 \] \[x+6y=12 \] so solved for x and got x=12-6y ?

OpenStudy (anonymous):

yes, that sounds good.

OpenStudy (anonymous):

Ok, now if you multiply the first equation by 15, you'll get 5x + 3y = 15

OpenStudy (anonymous):

then i would have to type in for x/3 as 12-6y/3 +y/5 = 1 ?

OpenStudy (anonymous):

so now if I take you first equation ninab and substitute into Cliff's I get: 5 (12 - 6 y) + 3y = 15 Can you take it from there?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

60-30y+3y=15 60-27y=15 ?

OpenStudy (anonymous):

yes, looks right so far

OpenStudy (anonymous):

so then subtract 15 from both sides i get 45=27y

OpenStudy (anonymous):

so y=5/3

OpenStudy (anonymous):

That seems right. So what about x?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

thank you both @CliffSedge and @Hugo47

OpenStudy (anonymous):

So since you had two equations, your solution is an ordered pair in the form (x,y) So (2, 5/3).